# Time delay interferometry with minimal null frequencies

### v1

@@ Abstract
Time delay interferometry (TDI) is a key technique employed in gravitational wave (GW) space missions to mitigate laser frequency noise by combining multiple laser links and establishing an equivalent equal arm interferometry. The null frequencies will be introduced in noise spectra and GW response when the periodical signal/noise is canceled in synthesized laser links. These frequencies are characteristic frequencies (CFs) of a TDI which related to its geometry of combination. In this work, we propose a second-generation TDI configuration referred to as hybrid Relay, whose CFs are only one-quarter that of the fiducial second-generation Michelson observables. We examine the performance of novel TDI configuration in laser noise cancellation and clock noise suppression and justify its essential capabilities. To assess its robustness for signal extraction, we simulate data containing GW signals from massive black hole binaries and perform parameter inferences with comparisons against the fiducial Michelson TDI configuration. The results demonstrate that the newly proposed TDI solution could be more robust than Michelson in fulfilling data analysis.
@@ I Introduction
In this study, we propose an alternative TDI observables referred to as hybrid Relay. These observables have only one-quarter CFs of second-generation Michelson, resulting in smoother noise spectra and GW response functions. As a second-generation TDI configuration, the new observables could sufficiently suppress laser frequency noise and clock noise as the Michelson observables. Another advantage is that the optimal TDI channels from the new configuration would be more robust than Michelson, avoiding susceptibility to unequal arm lengths during orbital motion. To validate the performance of the proposed TDI configuration, simulation and analysis are conducted using SATDI [29]. With the data containing GW signals from massive binary black holes (MBBH), the hybrid Relay demonstrates a stable capability to yield reliable results, making it a promising TDI candidate for space missions GW data processing.
@@ II Time delay interferometry of hybrid Relay
More lucid expressions of three TDI observables are illustrated through geometric diagrams in Fig. 1 [31, 32]. The vertical lines represent the trajectories of three S/C ($⃝i$ (), and duplicated S/C2 and S/C3 are plotted in the Relay-U and - diagrams to prevent intersections around non-integer delay ticks. The blue and magenta lines depict the combined links for the virtual interferometer of two beams. The represents the mismatches for the two beams due to the orbital dynamics.
@@
and its diagram is shown in the left plot of Fig. 2. The Y1 and Z1 could be obtained through cyclical permutation of links’ indexes. Hybrid Relay observables could be generated by combining U and delayed , denoted as U, and vice versa. They can be expressed as
@@ III.1 Suppressions of laser noise and clock noise
From Fig. 3, we can notice that CFs of X1 for either laser noise, secondary noise, or (original) clock noise emerge at and is the arm length of constellation. In contrast, the CFs of U only appear at the Hz. With only one-fourth of CFs, the noise spectra of hybrid Relay are smoother and more stable than those of Michelson On the other side, the frequencies imply that the integral circle of noise/signal in a single inter-S/C links representing the minimum common CFs that all TDI observables will possess. And hybrid Relay observables (U, V, W) only have these minimum CFs.
@@ III.2 Average sensitivities
The sensitivities of fiducial Michelson TDI channels are also calculated and depicted using dashed curves for comparison. The sensitivities of A and A are identical, and the E channels from different TDI configurations are also expected to be equal. However, the sensitivities of T differ from the science channels. In the frequency band higher than 50 mHz, the T could be slightly more sensitive than other observables. Due to the unequal arm, T diverges from T at 15 mHz and becomes as sensitive as the A/E for frequencies lower than 0.3 mHz, as analyzed in [40].
@@ IV.1 Data simulation
The data simulation is performed by using SATDI [29]. Noise data are generated with a sampling rate of 4 Hz for 30 days, and TDI is implemented by using interpolation [41]. The noise sources include acceleration noises and optical metrology noise by assuming that laser noise and clock noise are sufficiently suppressed. Simulations are curried out for second-generation Michelson (X1, Y1, Z1) and hybrid Relay (U, V, W). Optimal observables are generated by using Eq. (6) and employed to evaluate their analytical capabilities.
@@
The waveforms of source2 approaching the merger are depicted in Fig. 5. The x-axis is the barycentric dynamical time (TDB) with respect to the merger at the solar-system barycentric (SSB) center. The grey curve illustrates the GW signal arriving at the SSB center, and the blue and orange curves represent the responded waveforms in A and A observables, respectively. As displayed in the plot, the amplitudes of waveforms are modulated by the TDI response functions. In A, the signal amplitude is significantly suppressed around its CFs. Benefiting from the minimal CFs, the waveform in A is less restrained. Additionally, phase aliasing occurs at high frequencies due to the longer time spans of TDI compared to the GW periods. Frequency-domain waveforms are obtained via Fourier transform and presented in the lower panel of Fig. 5. The noise PSDs of A and A are plotted for comparison, revealing clearer suppression of waveforms at their CFs. In the Michelson observable A, the waveform amplitude quickly drops at each CF, as does its PSD. The waveform in A channel is subjected to a only few CFs.
@@ IV.2 Parameter inferences
The parameters of injected binaries are estimated by using Bayesian algorithm. The inferences are conducted utilizing the optimal observables from second-generation hybrid Relay (A, E, T) and Michelson (A, E, T) to compare their performances. The likelihood function of parameter estimation is given by [45]
@@
The inferred distributions for parameters of source1 are shown in Fig. 6, and results from three TDI combinations are presented. The green colors represent the inferred parameter distributions from (A, E, T), the blue hues signify estimation from (A, E, T), and the magentas depict the result of (A, E). Among these three combinations, with the exception of (A, E, T), the inferences from other two combinations are consistent, and injected parameters fall within sensible ranges. The discrepancy between (A, E) and (A, E, T) suggests that the T undermines the parameter estimation.
The inferred distributions for source2 are illustrated in Fig. 7. The result from the dataset (A, E, T) noticeably deviates from the injected values. After the T data is removed, the science data streams (A, E) yield a reasonable result that agrees with the distributions from (A, E, T). This further confirms that T data is problematic for signal analysis. Additionally, we can notice that inferred mass parameters () are (almost) beyond the creditable regions for the (A, E, T) group, and this could be caused by the discrepancies between the time-domain waveform model and frequency-domain waveform approximant [50]. For more massive sources1 binary, its waveform has fewer cycles than source2 over the span of 25 days. As a result, the current faithfulness between time-domain waveform and frequency-domain waveform models may allow for reasonable estimation of the mass parameters for source1, but might not meet the matching requirements for estimating the masses of source2. In the results from two sources, the hybrid Relay demonstrates robust outcomes with its full optimal data set, while the Michelson configuration is susceptible to bias introduced by its T channel. When comparing the results of Michelson across two sources, it appears that source1 exhibits less bias than source2, possibly because the A and E exert greater contributions than the T channel in the low-frequency band.
@@ IV.3 Post-analysis
By employing the waveform of source2, the values of for the optimal TDI channels are shown in Fig. 8. The upper left panel shows the curves for (A, E, T), and the upper right plot presents the results for (A, E, T). Grey vertical lines mark the CFs of each TDI configuration. The curves illustrate the science channels, A and E, dominate the SNR in the lower frequencies. The T makes a trivial contribution in the low frequency band, while the T exhibits a moderate SNR contribution in the same range. Our previous analysis showed that this gain is produced by the difference between the ordinary channels (X1, Y1, Z1). In other words, the signal in T is from the imperfect cancellation of X1, Y1, and Z1 [40]. That means T is susceptible to variations in interferometer arms over time. As a consequence, the T channel from the Michelson configuration not only faces intrinsic larger uncertainty from changing arm lengths due to orbital motion but also has instability in its response function to GW. In this analysis, PSDs are estimated by averaging data over 30 days, leading to inaccurate estimations introduced in T. As a result, including the T data stream biased the inference results.
@@
For the Michelson channels, another potential problem is the slight difference in CFs between the science channels, A and E. The PSD ratios of A and E, , for Michelson and hybrid Relay, are depicted in Fig. 9. Below 50 mHz, the PSD of A and E are closely equal, but as the frequency increase, their PSDs diverge especially around their CFs. In this data chunk, the CF of A is slightly lower than E which causes the ratio jumps from low to high at each CF. The rhythm of CFs in the hybrid Relay is more stable, the two channels have the same CFs. The ratios are dropped at each CF due to the PSD of A is lower than the E’s in current case. The asynchronous CFs in Michelson result from the different unequal arms involved in the channels. Optimal observables from the hybrid Relay have the same CFs, potentially enhancing the stability of their covariance matrix and thus benefiting the robustness of data analysis.
@@ V Conclusion and discussion
Machine learning techniques have been employed to boost the efficiency of GW identifications and estimations [54, 55, 56, 57, 58, 59, 60, 61, 62, and reference therein]. The concept of global fits has been proposed to simultaneously resolve various GW signals [63, 64]. TDI observables with stable PSDs and GW response functions are essential for these analyses. The proposed hybrid Relay could be the eligible configuration to fulfill these requirements, and we are committed to further studies in future work.

### v2

@@ Abstract
Time delay interferometry (TDI) is a key technique employed in gravitational wave (GW) space missions to mitigate laser frequency noise by combining multiple laser links and establishing an equivalent equal arm interferometry. The null frequencies will be introduced in noise spectra and GW response when the periodical signal/noise is canceled in synthesized laser links. These frequencies are characteristic frequencies (CFs) of a TDI which related to its geometry of combination. In this work, we implement a second-generation TDI configuration referred to as hybrid Relay to perform noise suppressions and data analysis, whose CFs are only one-quarter that of the fiducial second-generation Michelson observables. We examine the performance of TDI configuration in laser noise cancellation and clock noise suppression and justify its essential capabilities. To assess its robustness for signal extraction, we simulate data containing GW signals from massive black hole binaries and perform parameter inferences with comparisons against the fiducial Michelson TDI configuration. The results demonstrate that the alternative TDI solution could be more robust than Michelson in fulfilling data analysis.
@@ I Introduction
In this study, we implement an alternative TDI configuration referred to as hybrid Relay to perform noise suppressions and data analysis. Wang [29] first developed this TDI configuration in 2011 and analyzed its average sensitivity in 2020 [30]. Muratore et al. [31] also identified this configuration in 2020. One merit of these observables is that they have only one-quarter CFs of second-generation Michelson, resulting in smoother noise spectra and GW response functions. As a second-generation TDI configuration, the new observables could sufficiently suppress laser frequency noise and clock noise as the Michelson observables. Another advantage is that the optimal TDI channels from the alternative configuration would be more robust than Michelson, avoiding susceptibility to unequal arm lengths during orbital motion. To validate the performance of the proposed TDI configuration, simulation and analysis are conducted using SATDI [32]. With the data containing GW signals from massive binary black holes (MBBH), the hybrid Relay demonstrates a stable capability to yield reliable results. Further investigation shows that the hybrid Relay also exhibits a more robust ability to characterize the instrumental noises compared to the Michelson configuration [33]. These advantages make it a promising TDI candidate for space missions GW data processing.
@@ II Time delay interferometry of hybrid Relay
More lucid expressions of three TDI observables are illustrated through geometric diagrams in Fig. 1 [29, 30]. The vertical lines represent the trajectories of three S/C ($i$⃝ (), and duplicated S/C2 and S/C3 are plotted in the Relay-U and - diagrams to prevent intersections around non-integer delay ticks. The blue and magenta lines depict the combined links for the virtual interferometer of two beams. The represents the mismatches for the two beams due to the orbital dynamics.
@@
and its diagram is shown in the left plot of Fig. 2. The Y1 and Z1 could be obtained through cyclical permutation of links’ indexes. The first hybrid Relay observable could be generated by combining U and delayed , denoted as U [29, 30], and vice versa. They can be expressed as
@@
The hybrid Relay is chosen to minimize null frequencies as a second-generation TDI configuration and is applied to reduce clock jitter with current method. Second-generation TDI channels are typically formed by combining two identical first-generation TDI channels with a time shift [12, 18]. These configurations introduce null frequencies at , where , is the arm length, and is the relative time shift between two combined TDI paths. For instance, the time shift in X1 is as exemplified in Eq. (3), resulting in null frequencies at , as shown in next section. More second-generation TDI observables could be composed by synthesizing first-generation TDI channels at different S/Cs as developed in [35], and the interferometric beams are assembled with multiple beams. These observables may require a modified method for clock-jitter noise reduction, as current clock noise mitigation is applicable for two continuous beams [27]. The hybrid Relay is combined with two distinct Relay channels, which can counteract their path mismatches, thus avoiding additional null frequencies due to time shift. Additionally, it can reduce clock noise using the current algorithm without any modification, as demonstrated in following section.
@@ III.1 Suppressions of laser noise and clock noise
From Fig. 3, we can notice that CFs of X1 for either laser noise, secondary noise, or (original) clock noise emerge at and is the arm length of constellation. In contrast, the CFs of U only appear at the Hz. With only one-fourth of CFs, the noise spectra of hybrid Relay are smoother and more stable than those of Michelson On the other side, the frequencies imply that the noise/signal cancellation occurs with a factor of resulting from by a relative time delay of multiple . Except the ordinary channels of first-generation TDI Sagnac, this factor commonly included in the noise PSD and GW response function of TDI channels, and we deduce that would be the minimum CFs for these TDI observables. And hybrid Relay observables (U, V, W) only have these minimum CFs.
@@ III.2 Average sensitivities
The sensitivities of fiducial Michelson TDI channels are also calculated and depicted using dashed curves for comparison. The sensitivities of A and AX1 are identical, and the sensitivities of E channels from two TDI configurations are consistent with their A channels. However, the sensitivities of T differ from the science channels. In the frequency band higher than 50 mHz, the T could be slightly more sensitive than other observables. Due to the unequal arm, TX1 diverges from T at 15 mHz and becomes as sensitive as the A/E for frequencies lower than 0.3 mHz, as analyzed in [42, 43].
@@ IV.1 Data simulation
The data simulation is performed by using SATDI [32]. Noise data are generated with a sampling rate of 4 Hz for 30 days, and TDI is implemented by using interpolation [44]. The noise sources include acceleration noises and optical metrology noise by assuming that laser noise and clock noise are sufficiently suppressed. Simulations are carried out for second-generation Michelson (X1, Y1, Z1) and hybrid Relay (U, V, W). Optimal observables are generated by using Eq. (6) and employed to evaluate their analytical capabilities.
@@
The waveforms of source2 approaching the merger are depicted in Fig. 5. The x-axis is the barycentric dynamical time (TDB) with respect to the merger at the solar-system barycentric (SSB) center. The grey curve illustrates the GW signal arriving at the SSB center, and the blue and orange curves represent the responded waveforms in AX1 and A observables, respectively. As displayed in the plot, the amplitudes of waveforms are modulated by the TDI response functions. In AX1, the signal amplitude is significantly suppressed around its CFs. Benefiting from the minimal CFs, the waveform in A is less restrained. Additionally, phase aliasing occurs at high frequencies due to the longer time spans of TDI compared to the GW periods. Frequency-domain waveforms are obtained via Fourier transform and presented in the lower panel of Fig. 5. The noise PSDs of AX1 and A are plotted for comparison, revealing clearer suppression of waveforms at their CFs. In the Michelson observable AX1, the waveform amplitude quickly drops at each CF, as does its PSD. The waveform in A channel is subjected to a only few CFs.
@@ IV.2 Parameter inferences
The parameters of injected binaries are estimated by using Bayesian algorithm. The inferences are conducted utilizing the optimal observables from second-generation hybrid Relay (A, E, T) and Michelson (AX1, EX1, TX1) to compare their performances. The likelihood function of parameter estimation is given by [48]
@@
The inferred distributions for parameters of source1 are shown in Fig. 6, and results from three TDI combinations are presented. The green colors represent the inferred parameter distributions from (A, E, T), the blue hues signify estimation from (AX1, EX1, TX1), and the magentas depict the result of (AX1, EX1). Among these three combinations, with the exception of (AX1, EX1, TX1), the inferences from other two combinations are consistent, and injected parameters fall within sensible ranges. The discrepancy between (AX1, EX1) and (AX1, EX1, TX1) suggests that the TX1 undermines the parameter estimation.
The inferred distributions for source2 are illustrated in Fig. 7. The result from the dataset (AX1, EX1, TX1) noticeably deviates from the injected values. After the TX1 data is removed, the science data streams (AX1, EX1) yield a reasonable result that agrees with the distributions from (A, E, T). This further confirms that TX1 data is problematic for signal analysis. Additionally, we can notice that inferred mass parameters () are (almost) beyond the creditable regions for the (A, E, T) group, and this could be caused by the discrepancies between the time-domain waveform model and frequency-domain waveform approximant [53]. For more massive sources1 binary, its waveform has fewer cycles than source2 over the span of 25 days. As a result, the current faithfulness between time-domain waveform and frequency-domain waveform models may allow for reasonable estimation of the mass parameters for source1, but might not meet the matching requirements for estimating the masses of source2. In the results from two sources, the hybrid Relay demonstrates robust outcomes with its full optimal data set, while the Michelson configuration is susceptible to bias introduced by its T channel. When comparing the results of Michelson across two sources, it appears that source1 exhibits less bias than source2, possibly because the A and E exert greater contributions than the T channel in the low-frequency band.
@@ IV.3 Post-analysis
By employing the waveform of source2, the values of for the optimal TDI channels are shown in Fig. 8. The upper left panel shows the curves for (A, E, T), and the upper right plot presents the results for (AX1, EX1, TX1). Grey vertical lines mark the CFs of each TDI configuration. The curves illustrate the science channels, A and E, dominate the SNR in the lower frequencies. The T makes a trivial contribution in the low frequency band, while the TX1 exhibits a moderate SNR contribution in the same range. Our previous analysis showed that this gain is produced by the difference between the ordinary channels (X1, Y1, Z1). In other words, the signal in TX1 arises from the imperfect cancellation of X1, Y1, and Z1 [43], rendering it susceptible to variations in interferometer arms over time. And this residual signal in T channel is tricky to be accurately modeled in frequency-domain. Consequently, inaccurate signal subtraction in the TX1 channel could introduce bias into the analysis. In this analysis, PSDs are estimated by averaging data over 30 days, which may also leading to inaccurate estimations in TX1. As a result, including the TX1 data stream biased the inference results.
@@
For the Michelson channels, another potential problem is the slight difference in CFs between the science channels, A and E. The PSD ratios of A and E, , for Michelson and hybrid Relay, are depicted in Fig. 9. Below 50 mHz, the PSD of AX1 and EX1 are closely equal, but as the frequency increase, their PSDs diverge especially around their CFs. In this data chunk, the CF of AX1 is slightly lower than EX1 which causes the ratio jumps from low to high at each CF. The rhythm of CFs in the hybrid Relay is more stable, the two channels have the same CFs. The ratios are dropped at each CF due to the PSD of A is lower than the E’s in current case. The asynchronous CFs in Michelson result from the different unequal arms involved in the channels. Optimal observables from the hybrid Relay have the same CFs, potentially enhancing the stability of their covariance matrix and thus benefiting the robustness of data analysis.
@@ V Conclusion and discussion
Machine learning techniques have been employed to boost the efficiency of GW identifications and estimations [57, 58, 59, 60, 61, 62, 63, 64, 65, and reference therein]. The concept of global fits has been proposed to simultaneously resolve various GW signals [66, 67]. TDI observables with stable PSDs and GW response functions are essential for these analyses. The proposed hybrid Relay could be the eligible configuration to fulfill these requirements, and we are committed to further studies in future work.